### v1

@@ Abstract
In the simulations, we consider a range of total data block lengths $N=N_{\rm Z}+N_{\rm X}$ from $3000$ to $12000$ bits, and take $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the probability that each of Alice and Bob selects the $Z$ ($X$) basis. Also, we set $\omega=10^{-4}$ due to the finite extinction ratio of the intensity modulator [29]. The intensities $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the secret key rate per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests, the average behaviour of a typical fiber and detector model is considered (see Appendix B for further details). In addition, we assume an overall system loss of 30 dB, a dark count probability of Bob’s detectors of $p_{\rm d}=6\times{}10^{-7}$ [31] and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [19]. The EC leakage is set as in the BBM92 protocol, i.e. $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$. Note, however, that $\theta^{\rm th}$ here denotes the correctable QBER of the key block —which is different from the threshold single-photon error rate, $p^{\rm th}$, in the BBM92 scheme— and we set it to the average QBER of the considered channel model for consistency. Lastly, we set again the overall failure probability bound to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [32]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis. Overall, this leads to $\epsilon_{\rm sec}\approx{}2\times{}10^{-7}$.
The results are shown in Fig. 2b. As seen in the figure, our approach can reduce the minimum block size for key generation by more than $22\%$ compared to previous approaches. The origin of this improvement is twofold: a tighter estimation of the decoy-state bounds (thanks to the use of the Neyman constructions from Proposition 1), and a sharper random sampling bound (provided by Proposition 2).
@@
Conclusions. The performance of a QKD system is strongly affected by the finite statistics of its security proof. For instance, in satellite-based QKD, the raw data collection per satellite overpass [6, 10, 11, 9] is inevitably restricted, and tight statistical bounds may result in an enormous saving of time and resources. Furthermore, one could envision future QKD applications requiring short keys on-demand for daily transactions with high security standards. For those, a poor statistical analysis could result in an unbearable penalty on the security guarantees.
Here, we have developed two simple analytical tools for the finite statistics of QKD protocols, ultimately relying on the long-known additive Chernoff bound [26]. The first tool is a random sampling bound well-suited for the vast family of security proofs based on the notion of phase-error rate. The second tool, which is a byproduct of the first one, is a pair of Neyman constructions for averages of Bernoulli variables that naturally fit in security proofs of e.g. decoy-state QKD protocols [19]. Explicit comparison with prior work shows that our tools provide a neat advantage in many different parameter regimes relevant to QKD, generally downsizing the necessary block sizes for key extraction. In fact, the proposed random sampling bound surpasses existing alternatives [22, 15] in a broad parameter regime alien to QKD as well, which suggests that it might also be of interest to other applications. After all, this bound could in principle be used in disciplines well beyond the realm of QKD, such as quality control tests, clinical trials and hypothesis testing in general.
Acknowledgements.
@@
We acknowledge support from the Galician Regional Government (consolidation of Research Units: AtlantTIC), the Spanish Ministry of Economy and Competitiveness (MINECO), the Fondo Europeo de Desarrollo Regional (FEDER) through the grant No. PID2020-118178RB-C21, MICIN with funding from the European Union NextGenerationEU (PRTR-C17.I1) and the Galician Regional Government with own funding through the “Planes Complementarios de I+D+I con las Comunidades Autónomas” in Quantum Communication, the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska-Curie Grant No. 101072637 (Project QSI) and the project “Quantum Security Networks Partnership” (QSNP, grant agreement No. 101114043). M.C. acknowledges support from a “Salvador de Madariaga” grant from the Spanish Ministry of Science, Innovation and Universities (grant No. PRX22/00192).
Appendix A Ideal BBM92 protocol
A.1 Protocol description
@@
    Parameter estimation (PE): They publicly reveal the test data and compute the quantum bit error rate (QBER). The protocol is aborted if the QBER exceeds a tolerated threshold value $p^{\rm th}$.
@@
    Error verification (EV): They perform an EV step based on 2-universal hashing, using tags of $\log(2/\epsilon_{\rm cor})$ bits at most. If the EV tags do not match, the protocol is aborted. Otherwise, they proceed to privacy amplification.
@@
    Privacy amplification (PA): The parties perform a PA step based on 2-universal hashing, each of them obtaining a final key of length
    $l=\left\lfloor(N-n)\left\{1-h\left[q^{\rm th}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)\right]\right\}-\lambda_{\rm EC}-\log\left(\frac{1}{2\epsilon_{\rm cor}\epsilon_{\rm PA}^{2}}\right)\right\rfloor$ (7)
    bits. Here, $h(x)$ is the binary Shannon entropy function, while the function $q^{\rm th}_{N,n,\epsilon_{\rm PE}}(x)$ depends on the random sampling tool used for the parameter estimation. These are included in the next section. On the other hand, $\lambda_{\rm EC}$ is the EC leakage, $\epsilon_{\rm cor}$ is the correctness parameter and $\epsilon_{\rm PA}$ is the PA error.
@@ A.2 Failure probability
A.2.1 Additive Chernoff bound (this work)
@@
with $\Gamma^{+}_{n,\epsilon_{\rm PE}}(x)$ given by Eq. (2) of the main text. This corresponds to the solid blue line illustrated in Fig. 2a in the main text.
A.2.2 Serfling and Hush & Scovel inequalities combined [22]
Using Lemma 1 from [22], a candidate $q^{\rm th}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)$ is given by the solution to the following minimization problem:
$\begin{array}[]{ll}\min\limits_{(\xi,q^{\rm th})}&q^{\rm th},\\ \text{such that}&q^{\rm th}-p^{\rm th}>\xi,\\ &N\left(p^{\rm th}+\xi\right)\in\mathbb{Z^{+}},\\ &(N-n)^{2}\left(q^{\rm th}-p^{\rm th}-\xi\right)>1,\\ &\begin{aligned} &\epsilon_{\rm PE}=\exp\left(-\frac{2Nn\xi^{2}}{N-n+1}\right)+\exp\left\{-2F_{N(p^{\rm th+\xi})}\left[(N-n)^{2}\left(q^{\rm th}-p^{\rm th}-\xi\right)^{2}-1\right]\right\},\\ &\end{aligned}\end{array}$ (9)
where
$F_{N(p^{\rm th+\xi})}:=\frac{1}{N\left(p^{\rm th}+\xi\right)+1}+\frac{1}{N-N\left(p^{\rm th}+\xi\right)+1}.$ (10)
As explained in the Introduction of the main text, this approach provides an optimized combination of Serfling inequality [20] with Hush and Scovel inequality [23]. This corresponds to the dashed orange line plotted in Fig. 2a in the main text.
A.2.3 Serfling inequality [20]
Using Lemma 6 from [15] (which follows solely from Serfling inequality [20]), we have that
@@ A.2.3 Serfling inequality [20]
Indeed, one can easily show that the program of Eq. (9) can only yield a threshold equal or lower than Eq. (11). This latter equation corresponds to the dotted green line illustrated in Fig. 2a in the main text.
A.3 Simulations
Here we plot the secret key rate versus the block length for the ideal BBM92 protocol comparing the three methods above. Unlike Fig. 2a in the main text, which only considers block lengths up to $8000$, this figure goes up to $N=10^{6}$ bits and shows that the advantage of our method remains even in the large block size regime.
Figure 3: Finite secret key rate per raw data bit as a function of the data block size for the ideal BBM92 protocol with data block sizes up to $10^{6}$ bits. The solid blue line corresponds to Proposition 3 of the main text, while the dashed orange (dotted green) line refers to [22] ([15]).
Appendix B Decoy-state BB84 protocol
@@
B.1 Protocol description
Alice and Bob agree on the protocol inputs $T$, $q_{\rm X}$, $\mathcal{K}:=\left\{\mu,\nu,\omega\right\}$, $p_{\mu}$ $p_{\nu}$, $N_{\rm X}$, $N_{\rm Z}$, $\theta^{\rm th}$, $\lambda_{\rm EC}$, $\epsilon_{\rm PE}$, $\epsilon_{\rm cor}$, $\epsilon_{\rm PA}$, and $\delta$, specified in Table 2. In addition, if the random sampling tool of Proposition 2 in the main text (or Serfling inequality) is considered, they further agree on two more protocol inputs inherent to the corresponding PE test, $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$. Alternatively, if the tool of [22] is considered, the additional protocol inputs are $n_{1,\rm Z}^{\rm th,U(L)}$, $n_{1,\rm X}^{\rm th,U(L)}$ and $m_{1,\rm X}^{\rm th,U}$, with which they can ultimately define the relevant threshold values entering the secret key length formula:
@@ B.1 Protocol description
where $e_{1,\rm X}^{\rm th}=m_{1,\rm X}^{\rm th,U}/n_{1,\rm X}^{\rm th,L}$, =$\mathcal{B}=\bigl{\{}u\in\bigl{[}n_{1,\rm Z}^{\rm th,L}+n_{1,\rm X}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}+n_{1,\rm X}^{\rm th,U}\bigr{]},v\in\bigl{[}n_{1,\rm X}^{\rm th,L},n_{1,\rm X}^{\rm th,U}\bigr{]},u>v\bigr{\}},$ and the threshold function under consideration is that of Eq. (9). Importantly, we adopt a more contrived PE test when considering the random sampling tool of [22] in order to simplify the task of bounding the failure probability in Appendix B.3.2 below. Indeed, it is unclear whether a bound can be derived with this tool without modifying the PE test at all. After all, the tool in [22] is devised for the BBM92 protocol specifically, rather than for a decoy-state BB84 protocol, and in this work we provide a possible adaptation to the latter.
@@
    Sifting: The bases and intensity settings are publicly revealed and Alice and Bob identify the sets $\mathcal{Z}=\bigl{\{}i:a_{i}=b_{i}=Z,\ y_{i}^{\prime}\neq\emptyset\bigr{\}}$ and $\mathcal{X}=\left\{i:a_{i}=b_{i}=X,\ y_{i}^{\prime}\neq\emptyset\right\}$. $N_{\rm Z}$ rounds are drawn at random from $\mathcal{Z}$ to form the sifted key data, $\mathcal{Z}^{\rm sift}$, which decomposes as $\mathcal{Z}^{\rm sift}=\cup_{k\in\mathcal{K}}\mathcal{Z}^{\rm sift}_{k}$ with $\mathcal{Z}^{\rm sift}_{k}=\left\{i\in\mathcal{Z}^{\rm sift}:k_{i}=k\right\}$ of size $n_{{\rm Z},k}$. $N_{\rm X}$ rounds are drawn at random from $\mathcal{X}$ to form the test data, $\mathcal{X}^{\rm test}$, which decomposes as $\mathcal{X}^{\rm test}=\cup_{k\in\mathcal{K}}\mathcal{X}^{\rm test}_{k}$ with $\mathcal{X}^{\rm test}_{k}=\left\{i\in\mathcal{X}^{\rm test}:k_{i}=k\right\}$ of size $n_{{\rm X},k}$.
@@
    Parameter estimation (PE): For each $k\in\mathcal{K}$, they disclose the bit values in $\mathcal{X}_{k}^{\rm test}$ and compute the corresponding numbers of bit errors, $m_{{\rm X},k}$. With the available data, they perform the PE test that matches their preferred random sampling tool (see Appendix  B.3 for the details). If the test fails, they abort the protocol.
@@
    Error verification (EV): They perform an EV step based on 2-universal hashing, using tags of $\log(2/\epsilon_{\rm cor})$ bits at most. If the EV tags do not match, the protocol is aborted. Otherwise, they proceed to privacy amplification.
@@
    $l=\left\lfloor n^{\rm th}_{1,\mathrm{Z}}\left[1-h\left(\phi^{\rm th}_{1,\rm Z}\right)\right]-\lambda_{\rm EC}-\log\left(\frac{1}{2\epsilon_{\rm cor}\epsilon_{\rm PA}^{2}\delta}\right)\right\rfloor,$ (13)
which is provably $\left[2\left(\sqrt{\epsilon_{\rm PE}}+\delta\right)+\epsilon_{\rm PA}\right]$-secret and $\epsilon_{\rm cor}$-correct for the requested failure probability bound, $\epsilon_{\rm PE}$ [34, 33], carefully addressed in Appendix B.3.
@@
B.2 Decoy-state analysis
Here, we keep the analytical decoy-state bounds provided in [19] and generalize the statistical analysis, assuming a common error probability $\varepsilon$ per concentration inequality for simplicity.

### v2

@@ Abstract
Let us now assess the more practical decoy-state BB84 protocol. We consider a scheme with three common intensities per basis, $\mu$ (signal), $\nu$ (decoy) and $\omega$ (vacuum), and employ the $Z$ ($X$) basis for key extraction (PE) with a fixed-size data block of $N_{\rm Z}$ ($N_{\rm X}$) bits drawn at random from the corresponding data pool. Again, in the simulations we contemplate blocks up to $N=N_{\rm Z}+N_{\rm X}=10^{6}$ bits, and superpose an inset of the minimum block-size regime. Naturally, we set $N_{\rm Z}=q_{\rm Z}^{2}/(q_{\rm Z}^{2}+q_{\rm X}^{2})N$, where $q_{\rm Z}$ ($q_{\rm X}$) denotes the $Z$ ($X$) basis probability, common to both parties. Also, while we fix $\omega=10^{-4}$ to account for the finite extinction ratio of the intensity modulator [35], the settings $\mu$ and $\nu$, together with their respective probabilities, $p_{\mu}$ and $p_{\nu}$, and the test basis probability, $q_{\rm X}$, are tuned to optimize the key length per block, $l/N$. For simplicity, in order to select the acceptance thresholds of the PE tests [36], the average behaviour of a typical fiber and detector model is considered (see Supplemental Material II for further details). We assume an overall system loss of 30 dB, a dark count probability of $p_{\rm d}=6\times{}10^{-7}$ [37], and a misalignment error rate of $e_{\rm mis}=5\times 10^{-3}$ [12]. As in the BBM92 protocol, the EC leakage is set to $\lambda_{\rm EC}=1.19N_{\rm Z}h(\theta^{\rm th})$, $\theta^{\rm th}$ denoting the correctable QBER of a key block [38]. Lastly, we set again the overall failure probability to $\epsilon_{\rm PE}=4\times 10^{-16}$, and assume a common value for each individual error term contributing to it [39]. Similarly, we take $\epsilon_{\rm cor}=\epsilon_{\rm PA}=\delta=10^{-8}$, where $\delta>0$ is an intrinsic variable of the security analysis (see Supplemental Material II). Overall, this leads to $\epsilon_{\rm sec}={}7\times{}10^{-8}$.
The results are shown in Fig. 2b. Unlike the BBM92 protocol, the finite-key security of the decoy-state scheme does not rely on the random sampling tool alone, but also on CIs for the average parameter of a set of Bernoulli trials. Indeed, such CIs are repeatedly used in the decoy-state analysis, and thus they have a very critical impact in the performance. As an example, both orange lines in Fig. 2b use the random sampling tool of [15], but the dotted line relies on Hoeffding inequality [17] for the extra CIs (as originally done in [12]), while the dashed one relies on the multiplicative Chernoff bound [20] (following the improvement of [18]). Meanwhile, the solid blue line results from using the random sampling tool of Proposition 2 and the extra CIs of Proposition 1, both originating from our relaxed additive Chernoff bound and each providing an advantage on its own. Lastly, the solid red line uses our numerical upgrades of these two propositions, deploying optimal CIs in both cases. Overall, our analytical tools reduce the minimum block size by more than $22\%$ compared to previous approaches, and the reduction exceeds $38\%$ with the numerical tools. Once more, our analytical and numerical solutions perform very similarly in non-critical regimes, with a strong edge over existing approaches.
@@
Conclusions. The performance of QKD is deeply affected by finite statistics. In satellite QKD [34, 40, 41, 42], the data collection per satellite overpass is limited, and sharp statistics may translate into an enormous saving of time and resources. What is more, one could envision future QKD applications requiring short keys on-demand for daily transactions with high security standards. For these, a poor statistical analysis could result in an unbearable penalty.
In this work, we have developed simple analytical tools of unprecedented tightness among EUR-based QKD security proofs [5]. Technically, the key enablers of these tools are a careful relaxation of the long-known additive Chernoff bound [20], and a classical link between random sampling with and without replacement [17]. Arguably, the tools might be of interest beyond the realm of QKD, in applications such as quality control tests, clinical trials or hypothesis testing in general. On top of it, we have put forward the possibility of replacing tail inequalities —the cornerstone of statistical analyses in QKD today— by exact computational methods in statistical inference. Remarkably, our results reveal the game-changing potential of such methods to achieve QKD with minimum block sizes, and reinforce the tightness of our simpler analytical tools when moderate-to-large blocks are considered.
@@
Acknowledgements. We acknowledge fruitful discussions with Jay Bartroff, Davide Rusca and Jean-Daniel Bancal. This work was supported by the Galician Regional Government (consolidation of Research Units: AtlantTIC), the Spanish Ministry of Economy and Competitiveness (MINECO), the Fondo Europeo de Desarrollo Regional (FEDER) through the grant No. PID2020-118178RB-C21, MICIN with funding from the European Union NextGenerationEU (PRTR-C17.I1) and the Galician Regional Government with own funding through the “Planes Complementarios de I+D+I con las Comunidades Autónomas” in Quantum Communication, the European Union’s Horizon Europe Framework Programme under the Marie Sklodowska-Curie Grant No. 101072637 (Project QSI) and the project QSNP (grant agreement No. 101114043). M.C. acknowledges support from a “Salvador de Madariaga” grant from the Spanish Ministry of Science, Innovation and Universities (grant No. PRX22/00192).
I Appendix I: ideal BBM92 protocol
I.1 Protocol description
@@
    Parameter estimation (PE): They publicly reveal the test data and compute the quantum bit error rate (QBER). If the QBER exceeds a tolerated threshold value $p^{\rm th}$, they abort the protocol.
@@
    Error correction (EC): They run a pre-defined EC protocol to reconcile their sifted keys, revealing $\lambda_{\rm EC}$ secret-key bits at most.
@@
    Error verification (EV): They perform an EV step based on 2-universal hashing, using tags of $\log(2/\epsilon_{\rm cor})$ bits at most. If the EV tags do not match, they abort the protocol. Otherwise, they proceed to privacy amplification.
@@
    Privacy amplification (PA): The parties perform a PA step based on 2-universal hashing, each of them obtaining a final key of $l$ bits, with
    $l=\left\lfloor(N-n)\left\{1-h\left[q^{\rm th}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)\right]\right\}-\lambda_{\rm EC}-\log\left(\frac{1}{2\epsilon_{\rm cor}\epsilon_{\rm PA}^{2}}\right)\right\rfloor.$ (11)
    Here, $h(x)$ is the binary entropy function, and the threshold function $q^{\rm th}_{N,n,\epsilon_{\rm PE}}(x)$ depends on the random sampling tool used for PE. These are included in the next section. On the other hand, $\lambda_{\rm EC}$ is the EC leakage, $\epsilon_{\rm cor}$ is the correctness parameter and $\epsilon_{\rm PA}$ is the PA error.
@@ I.2 Failure probability
I.2.1 Relaxed additive Chernoff bound
@@
where $\Gamma^{+}_{n,\epsilon_{\rm PE}}(x)$ is given in Proposition 1 of the main text. This corresponds to the solid blue line of Fig. 2a there.
I.2.2 Serfling and Hush & Scovel inequalities combined
From Lemma 1 in [15], we have
$q^{\rm th}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)=\max_{\xi}\left\{p^{\rm th}+\xi+\frac{1}{N-n}\sqrt{1+\frac{1}{2g}\ln\left[\frac{1}{\epsilon_{\rm PE}-\displaystyle{\exp\left(-\frac{2Nn\xi^{2}}{N-n+1}\right)}}\right]}\right\}$ (13)
with
$g=\frac{1}{N\left(p^{\rm th}+\xi\right)+1}+\frac{1}{N\left(1-p^{\rm th}-\xi\right)+1},$ (14)
where the maximization runs over all $\xi>0$ such that $N(p^{\rm th}+\xi)\in\mathbb{Z}^{+}$ and $\epsilon_{\rm PE}>\displaystyle{}{\exp\left(-\frac{2Nn\xi^{2}}{N-n+1}\right)}$.
As mentioned in the main text, this approach provides an optimized combination of Serfling inequality [14] with Hush and Scovel inequality [16], and it corresponds to the dashed orange line of Fig. 2a in the main text.
I.2.3 Serfling inequality
Using Lemma 6 from [8] (which follows from Serfling inequality [14]), we have
@@ I.2.3 Serfling inequality
This corresponds to the dotted green line of Fig. 2a in the main text.
I.2.4 Exact hypergeometric cumulative mass function
In the main text, a numerical confidence upper bound $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ on the population parameter is presented [27], based on the classical Clopper-Pearson method of pivoting the cumulative mass function [19]. Since one can replace $\Gamma^{+}_{n,\epsilon}(\hat{p})$ by $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ in Corollary 1 of the main text, $\mathcal{CP}^{+}_{N,n,\epsilon}(\hat{p})$ automatically fulfills Proposition 2 there as well. From this, one can easily show that it also satisfies Proposition 3, thereby providing another suitable threshold function:
$q^{\rm th}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)=\frac{N\mathcal{CP}^{+}_{N,n,\epsilon_{\rm PE}}\left(p^{\rm th}\right)-n{}p^{\rm th}}{N-n}.$ (16)
This corresponds to the solid red line in of Fig. 2a in the main text.
II Appendix II: decoy-state BB84 protocol
@@
II.1 Protocol description
Alice and Bob agree on the protocol inputs $T$, $q_{\rm X}$, $\mathcal{K}:=\left\{\mu,\nu,\omega\right\}$, $p_{\mu}$ $p_{\nu}$, $N_{\rm X}$, $N_{\rm Z}$, $\theta^{\rm th}$, $\lambda_{\rm EC}$, $\epsilon_{\rm PE}$, $\epsilon_{\rm cor}$, $\epsilon_{\rm PA}$, and $\delta$, specified in Table 2. In addition, unless the random sampling tool of [15] is considered, they further agree on two more protocol inputs inherent to the corresponding PE test, $n_{1,\rm Z}^{\rm th}$ and $\phi_{1,\rm Z}^{\rm th}$. Alternatively, if the tool of [15] is considered, the additional protocol inputs are $n_{1,\rm Z}^{\rm th,U(L)}$, $n_{1,\rm X}^{\rm th,U(L)}$ and $m_{1,\rm X}^{\rm th,U}$, with which they can ultimately define the relevant threshold values entering the secret key length formula:
@@ II.1 Protocol description
where $e_{1,\rm X}^{\rm th}=m_{1,\rm X}^{\rm th,U}/n_{1,\rm X}^{\rm th,L}$, =$\mathcal{B}=\bigl{\{}u\in\bigl{[}n_{1,\rm Z}^{\rm th,L}+n_{1,\rm X}^{\rm th,L},n_{1,\rm Z}^{\rm th,U}+n_{1,\rm X}^{\rm th,U}\bigr{]},v\in\bigl{[}n_{1,\rm X}^{\rm th,L},n_{1,\rm X}^{\rm th,U}\bigr{]},u>v\bigr{\}},$ and the threshold function under consideration is that of Eq. (13). Importantly, we adopt a finer-grained PE test for the random sampling tool of [15] in order to simplify the task of bounding the failure probability in Sec. II.3.2 below. Indeed, it is unclear whether a bound can be derived with this tool without modifying the PE test at all. After all, the tool in [15] is devised for the BBM92 protocol specifically, rather than for a decoy-state BB84 protocol, and in this work we provide a possible adaptation to the latter.
@@
    Sifting: The bases and intensity settings are publicly revealed and Alice and Bob identify the sets $\mathcal{Z}=\bigl{\{}i:a_{i}=b_{i}=Z,\ y_{i}^{\prime}\neq\emptyset\bigr{\}}$ and $\mathcal{X}=\left\{i:a_{i}=b_{i}=X,\ y_{i}^{\prime}\neq\emptyset\right\}$. $N_{\rm Z}$ rounds are drawn at random from $\mathcal{Z}$ to form the sifted key data, $\mathcal{Z}^{\rm sift}$, which decomposes as $\mathcal{Z}^{\rm sift}=\cup_{k\in\mathcal{K}}\mathcal{Z}^{\rm sift}_{k}$ with $\mathcal{Z}^{\rm sift}_{k}=\left\{i\in\mathcal{Z}^{\rm sift}:k_{i}=k\right\}$, of size $n_{{\rm Z},k}$. $N_{\rm X}$ rounds are drawn at random from $\mathcal{X}$ to form the test data, $\mathcal{X}^{\rm test}$, which decomposes as $\mathcal{X}^{\rm test}=\cup_{k\in\mathcal{K}}\mathcal{X}^{\rm test}_{k}$ with $\mathcal{X}^{\rm test}_{k}=\left\{i\in\mathcal{X}^{\rm test}:k_{i}=k\right\}$, of size $n_{{\rm X},k}$.
@@
    Parameter estimation (PE): For each $k\in\mathcal{K}$, they disclose the bit values in $\mathcal{X}_{k}^{\rm test}$ and compute the corresponding numbers of bit errors, $m_{{\rm X},k}$. With the available data, they perform the PE test that matches their preferred random sampling tool (see Sec. II.3 for the details). If the test fails, they abort the protocol.
@@
    Error verification (EV): They perform an EV step based on 2-universal hashing, using tags of $\log(2/\epsilon_{\rm cor})$ bits at most. If the EV tags do not match, they abort the protocol. Otherwise, they proceed to privacy amplification.
@@
    $l=\left\lfloor n^{\rm th}_{1,\mathrm{Z}}\left[1-h\left(\phi^{\rm th}_{1,\rm Z}\right)\right]-\lambda_{\rm EC}-\log\left(\frac{1}{2\epsilon_{\rm cor}\epsilon_{\rm PA}^{2}\delta}\right)\right\rfloor.$ (18)
This approach yields provably $\left[2\left(\sqrt{\epsilon_{\rm PE}}+\delta\right)+\epsilon_{\rm PA}\right]$-secret [44, 43] and $\epsilon_{\rm cor}$-correct keys for the requested failure probability bound, $\epsilon_{\rm PE}$, carefully addressed in Sec. II.3.
@@
II.2 Decoy-state analysis
Here, we keep the analytical decoy-state bounds provided in [12] and generalize the statistical analysis [46], assuming a common error probability $\varepsilon$ per statistical bound for simplicity.